Find the value of
1) 32⅕ (32 to the power of 1/5)
2) 32 ⅖ (32 to the power of 2/5)
Thanks :)

Found 3 solutions by sudhanshu_kmr, stanbon, RakshithC:
Answer by sudhanshu_kmr(1152)
Answer by stanbon(75887)

Find the value of
1) 32⅕ (32 to the power of 1/5)
---
The "5" in the denominator of the exponent means,
"take the 5th root".
Ans: 32^(1/5) = 2
--------------------------------------
2) 32 ⅖ (32 to the power of 2/5)
The "5" means 5th root.
The "2" means square.
--------
32^(2/5) = 2^2 = 4
=======================
Cheers,
Stan H.
=============

Answer by RakshithC(1)

32^(2/5) = , 5th root of 32 square. that means 5th root of 32*32. now 32*32= 1024, 5th root of 1024 = 4.
proof 4*4*4*4*4=1024.